OpenStudy (anonymous):

Solve the equation in the interval [0. 2pi) cosx sinx - sinx = 0

OpenStudy (anonymous):

sin(x) (cos(x)-1) = 0 => sin(x) = 0 OR cos(x) = 1 Find x from that...and in the interval given.

OpenStudy (anonymous):

The book says the answer is 0, pi but I don't understand how they got there

OpenStudy (anonymous):

sin(x) = 0 when x = 0 and x = 180 degrees (pi radians).

OpenStudy (anonymous):

cos(x) = 1 when x = 0 and x = 360 degrees, which is 2pi, but 2pi is NOT in the domain given. So, the right answer is 0 or pi.

OpenStudy (anonymous):

So question???? Did you just factor out the sinx out of this problem and set it equal to zero? And with cosx-1 left over you moved the 1 to the other side?

OpenStudy (anonymous):

YES. YES!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

When a product of two things is zero, either or both can be zero. So, when sin(x) (cos(x)-1) = 0, either sin(x) = 0 OR cos(x)-1 = 0. => sin(x) = 0 OR cos(x) = 1.